Passenger miles flown on Northeast​ Airlines, a commuter firm serving the Boston​ hub, are shown for the past 12​ weeks:

                                                                                                                                                                        

Week

1

2

3

4

5

6

7

8

9

10

11

12

Actual Passenger Miles​ (in thousands)

1515

1919

1717

2121

2020

1818

2222

2020

2424

2222

1212

2121

​a) Assuming an initial forecast for week 1 of

15,000

​miles, use exponential smoothing to compute miles for weeks 2 through 12. Use

alphaαequals=0.250.25

​(round your responses to two decimal​ places).

Week

1

2

3

4

5

6

7

8

9

10

11

12

Forecasted Passenger Miles​ (in thousands)

15.0015.00

15.0015.00

16.0016.00

16.2516.25

17.4417.44

18.0818.08

18.0618.06

19.0519.05

19.2919.29

20.4720.47

??

??
Answer #1

a) Using the exponential smoothing method

F(2) = 0.25*Actual(1) + (1-0.25)*F(1)

= 0.25*15 + 0.75*15

= 15

F(3) = 0.25*19 + 0.75*15

= 16

F(4) = 0.25*17 + 0.75*16

= 16.25

F(5) = 0.25*21 + 0.75*16.25

= 17.44

F(6) = 0.25*20 + 0.75*17.44

= 18.08

F(7) = 0.25*18 + 0.75*18.08

= 18.06

F(8) = 0.25*22 + 0.75*18.06

= 19.05

F(9) = 0.25*20 + 0.75*19.05

= 19.29

F(10) = 0.25*24 + 0.75*19.29

= 20.47

F(11) = 0.25*22 + 0.75*20.47

= 20.85

F(12) = 0.25*12 + 0.75*20.85

= 18.64